In a four digit number, the sum of the thousands and the hundreds digits is 3. The tens digit is four times the hundreds digit. The ones digit is seven more than the hundreds digit. no two digits are equal. What is the four digit number?
Answer by ankor@dixie-net.com(22740)

In a four digit number,
a, b, c, d
the sum of the thousands and the hundreds digits is 3.
a + b = 3
The tens digit is four times the hundreds digit.
c = 4b
The ones digit is seven more than the hundreds digit.
d = b + 7
no two digits are equal. What is the four digit number?
We know that b=1, a= 2 or b=2, a=1
if b=1, then a=2
c = 4(1)
c = 4
and
d = 1 + 7
d = 8
then 2148 is the number
:
If b=2 then a=1
c = 4(2)
c = 8
and
d = 2 + 7
c = 9
then 1289 is the number